1)

2)

)Periodic Inventory by Three Methods The units of an item available for sale during the year were as follows: Jan. 1 Inventory
Answer #1

1)

Date Particulars Units Unit cost Total cost
Jan. 1 Inventory 50 130 6500
Mar. 10 Purchases 60 138 8280
Aug. 30 Purchases 20 142 2840
Dec. 12 Purchases 70 144 10080
Goods available for sale 200 27700
Ending inventory units 80
FIFO method : In this method those goods are sold first, which are purchased first and the ending inventory is from recent purchases.
Merchandise inventory = ( 70 * 144 ) + ( 10 * 142 ) 11500
LIFO method : In this method those goods are sold first which are purchased last and the ending inventory is from beginning inventory and earlier purchases.
Merchandise inventory = ( 50 * 130 ) + ( 30 * 138 ) 10640
Weighted average cost method :
Weighted average cost per unit = Cost of goods available for sale / Unit available for sale = 27700 / 200 138.5
Merchandise inventory = 80 * 138.5 11080
Final answer :
Cost of merchandise inventory and cost of merchandise sold
Inventory method Merchandise inventory

Merchandise sold

[ Cost of goods available for sale - Merchandise inventory ]

First-in, first out (FIFO) 11500 16200
Last-in, First-out (LIFO) 10640 17060
Weighted average cost 11080 16620

2)

Date Particulars Units Unit cost Total cost
Jan. 1 Inventory 1100 120 132000
Feb. 17 Purchases 1405 122 171410
Jul. 21 Purchases 1560 125 195000
Nov. 23 Purchases 1130 125 141250
Goods available for sale 5195 639660
Enidng inventory units 1205
FIFO method : In this method those goods are sold first, which are purchased first and the ending inventory is from recent purchases.
Inventory cost = ( 1130 * 125 ) + ( 75 * 125 ) 150625
LIFO method : In this method those goods are sold first which are purchased last and the ending inventory is from beginning inventory and earlier purchases.
Inventory cost = ( 1100 * 120 ) + ( 105 * 122 ) 144810
Weighted average cost method :
Weighted average cost per unit = Cost of goods available for sale / Unit available for sale = 639660 / 5195 123.13
Inventory cost = 1205 * 123.13 148372